What is the mole fraction of a NaOH in an aqueous solution that contains 33.9% NaOH by mass?


1
Expert's answer
2021-03-04T06:33:35-0500

If the mass of solution is 100 g, than

m(NaOH)=33.9 g;

m(H2O)=100-33.9=66.1 g


n(NaOH)=m/Mr=33.9/40=0.8475 mol;

n(H2O)=m/Mr=66.1/18=3.67 mol;


Mole fraction of NaOH (XNaOH):

XNaOH= n(NaOH)/(n(NaOH)+n(H2O))*100%=0.8475/(0.8475+3.67)*100%=18.76%


Answer: mole fraction of NaOH is 18.76%